# Diffusion-based Extreme Image Compression with Compressed Feature Initialization

### v1

@@ 3.2 Accelerating Denoising Process with Relay Residual Diffusion
where we set $\lambda=\frac{\sqrt{\bar{\alpha}_{N}}}{\sqrt{1-\bar{\alpha}_{N}}}$ to ensures $\eta_{N}=1$. Detailed derivation is presented in Appendix A. Substituting Eq. (4) into Eq. (2), the diffusion process can be further written as follows:
$\bm{z}_{n}$ $=\sqrt{\bar{\alpha}_{n}}(\bm{z}_{0}+\lambda\frac{\sqrt{1-\bar{\alpha}_{n}}}{\sqrt{\bar{\alpha}_{n}}}\bm{e})+\sqrt{1-\bar{\alpha}_{n}}\epsilon_{n}$ (5)
$=\sqrt{\bar{\alpha}_{n}}\bm{z}_{0}+\sqrt{1-\bar{\alpha}_{n}}\underbrace{(\lambda\bm{e}+\epsilon_{n})}_{\tilde{\epsilon}_{n}}.$ (6)
Since Eq. (6) has the same structure as Eq. (1), we can easily incorporate stable diffusion into our framework. For the denoising process, the noise estimator $\epsilon_{\theta}$ is learned to predict $\tilde{\epsilon}_{n}$ at each time-step $n$. The optimization of noise estimator $\epsilon_{\theta}$ is defined as
$\mathcal{L}_{ne}$ $=\mathbb{E}_{\bm{z}_{0},\bm{z}_{c},\bm{c},n,\epsilon_{n}}\|\bm{z}_{0}-\hat{\bm{z}}_{0}\|^{2}_{2}$ (7)
$=\omega_{n}\mathbb{E}_{\bm{z}_{0},\bm{z}_{c},\bm{c},n,\epsilon_{n}}\|\tilde{\epsilon}_{n}-\epsilon_{\theta}(\bm{z}_{n},\bm{c},n)\|^{2}_{2},$ (8)
where $\omega_{n}=\frac{1-\bar{\alpha}_{n}}{\bar{\alpha}_{n}}$. After that, we can start from the compressed latent features $\bm{z}_{c}$ and reconstruct the image using Eq. 3 without knowing the residual $\bm{e}$.
3.3 Fixed-Step Fine-Tuning Strategy
Most existing diffusion-based image compression methods adopt the same training strategy as DDPM (Ho et al., 2020), where each time-step is trained independently. However, the lack of coordination among time-steps can lead to error accumulation and suboptimal reconstruction quality. To address this issue, we employ a two-stage training strategy. As shown in Fig. 3(a), we first train each time-step $n$ independently, allowing the model to learn to remove noise and residuals at each step. The optimization objective consists of the rate-distortion loss, codebook loss (Van Den Oord et al., 2017) and noise estimation loss:
$\mathcal{L}_{stage\ I}=\underbrace{\lambda_{r}\|\bm{z}_{0}-\bm{z}_{c}\|_{2}^{2}+R(\hat{\bm{y}})}_{rate-distortion\ loss\ \mathcal{L}_{rd}}+\underbrace{\|sg(\bm{l}_{p})-\hat{\bm{l}}_{p}\|_{2}^{2}+\beta\|sg(\hat{\bm{l}}_{p})-\bm{l}_{p}\|_{2}^{2}}_{codebook\ loss\ \mathcal{L}_{cb}}+\lambda_{r}\mathcal{L}_{ne},$ (9)
@@ 3.3 Fixed-Step Fine-Tuning Strategy
To this end, we further employ a fixed-step fine-tuning strategy to eliminate the discrepancy between the training and inference phases. As shown in Fig. 3(b), in each training step, we utilize spaced DDPM sampling (Nichol & Dhariwal, 2021) with $L$ fixed time-steps to reconstruct the noise-free latent features $\hat{\bm{z}}_{0}$ from the starting point $\bm{z_{N}}$ and map $\hat{\bm{z}}_{0}$ back to the pixel space $\hat{\bm{x}}=\mathcal{D}(\hat{\bm{z}}_{0})$. The loss function used in this stage is as follows:
$L_{stage\ II}=\mathcal{L}_{rd}+\mathcal{L}_{cb}+\lambda_{r}\|\bm{z}_{0}-\hat{\bm{z}}_{0}\|^{2}_{2}+\lambda_{r}(\|\bm{x}-\hat{\bm{x}}\|_{2}^{2}+\lambda_{lpips}\mathcal{L}_{lpips}(\bm{x},\hat{\bm{x}})),$ (10)
@@
Figure 3: The two-stage training strategy of RDEIC. (a) Independent training: we randomly pick a time-step $n$ and train each time-step $n$ independently. This ensures that the model effectively learns to remove added noise and residuals at every step. (b) Fixed-step fine-tuning: $L$ fixed denoising steps are used to iteratively reconstruct a noise-free latent features $\hat{\bm{z}}_{0}$ from $\bm{z}_{N}$, which is consistent with the inference phase.
3.4 Controllable Detail Generation
@@
Although the fixed-step fine-tuning strategy significantly improves reconstruction quality, it requires a fixed number of denoising steps in the inference phase, making it impossible to achieve a trade-off between smoothness and sharpness by adjusting the number of denoising steps (Li et al., 2024b). To address this limitation, we introduce a controllable detail generation method that allows us to dynamically balance smoothness and sharpness without being constrained by the fixed-step requirement, which enables more versatile and user-specific image reconstructions.
@@
Since the compressed latent feature already contains image information, directly using stable diffusion’s noise estimator $\epsilon_{sd}$ to predict noise $\epsilon_{sd}(\bm{z}_{n},n)$ results in low-frequency reconstructed images, as shown in the second column of Fig. 7 and Fig. 13. Inspired by classifier-free guidance (Ho & Salimans, 2021), we decompose the predicted noise $\epsilon_{\theta}(\bm{z}_{n},\bm{c},n)$ into a low-frequency component $\epsilon_{sd}(\bm{z}_{n},n)$ and a high-frequency component $\epsilon_{\theta}(\bm{z}_{n},\bm{c},n)-\epsilon_{sd}(\bm{z}_{n},n)$, and control the balance between smoothness and sharpness by adjusting the intensity of the high-frequency component:
$\hat{\epsilon}_{n}=\epsilon_{sd}(\bm{z}_{n},n)+\lambda_{s}(\epsilon_{\theta}(\bm{z}_{n},\bm{c},n)-\epsilon_{sd}(\bm{z}_{n},n)),$ (11)
@@
where $\lambda_{s}$ is the guidance scale. By adjusting the value of $\lambda_{s}$, we can regulate the amount of high-frequency details introduced into the reconstructed image.
4 Experiments
4.1 Experimental Setup
@@
Datasets. The proposed RDEIC is trained on the LSDIR (Li et al., 2023) dataset, which contains 84,911 high-quality images. For evaluation, we use three common benchmark datasets, i.e., the Kodak (Franzen, 1999) dataset with 24 natural images of 768$\times$512 pixels, the Tecnick (Asuni & Giachetti, 2014) dataset with 140 images of 1200$\times$1200 pixels, and the CLIC2020 (Toderici et al., 2020) dataset with 428 high-quality images. For the Tecnick and CLIC2020 datasets, we resize the images so that the shorter dimension is equal to 768 and then center-crop them with 768$\times$768 spatial resolution (Yang & Mandt, 2023).
@@
Implementation details. For training, we use the Adam (Kingma & Ba, 2014) optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$ for a total of 300K iterations. The batch size is set to 4. We use Stable Diffusion 2.1-base [^1] https://huggingface.co/stabilityai/stable-diffusion-2-1-base as the specific implementation of stable diffusion. Throughout all our experiments, the weights of stable diffusion remain frozen. To achieve different compression ratios, we train five models with $\lambda_{r}$ selected from {2, 1, 0.5, 0.25, 0.1}. The total number $N$ of denoising steps is set to 300. As described in Section 3.3, the training process is divided into two stages. 1) Independent training. During this stage, the initial learning rate is set to 1×$\times 10^{-4}$ and images are randomly cropped to 512$\times$512 patches. We first train the proposed RDEIC with $\lambda_{r}=2$ for 100K iterations. The learning rate is then reduced to 2×$\times 10^{-5}$ and the model is trained with target $\lambda_{r}$ for another 100K iterations. 2) Fixed-step fine-tuning. In this stage, the learning rate is set to 2×$\times 10^{-5}$ and images are randomly cropped to 256$\times$256 patches. We fine-tune the model through the entire reconstruction process for 100K iterations. When $\lambda_{r}\in\{2,1\}$, the fixed number $L$ is set to 2, otherwise, it is 5. All experiments are conducted on a single NVIDIA GeForce RTX 4090 GPU.
@@
Metrics. For quantitative evaluation, we employ several established metrics to measure the visual quality of the reconstructed images, including reference perceptual metrics LPIPS (Zhang et al., 2018), DISTS (Ding et al., 2020), FID (Heusel et al., 2017) and KID (Bińkowski et al., 2018) and no-reference perceptual metrics NIQE (Mittal et al., 2012). We also employ distortion metrics PSNR, SSIM and MS-SSIM (Wang et al., 2003) to measure the fidelity of reconstructions. Note that FID and KID are calculated on 256$\times$256 patches according to Mentzer et al. (2020).
@@
Comparison methods. We compare the proposed RDEIC with several representative extreme image compression methods, including the traditional compression standard: VVC (Bross et al., 2021); GANs-based compression methods: HiFiC (Mentzer et al., 2020), MS-ILLM (Muckley et al., 2023), and VQIR (Wei et al., 2024), and diffusion-based methods: Text+Sketch (Lei et al., 2023), PerCo (Careil et al., 2024),and DiffEIC (Li et al., 2024b). More details can be found in Appendix B.
Figure 4: Quantitative comparisons with state-of-the-art methods on the CLIC2020 dataset.
Table 1: BD-rate (%) for different methods on the CLIC2020 dataset.
Methods Perception Distortion Average
DISTS FID KID NIQE LPIPS PSNR MS-SSIM SSIM
HiFiC 241.7 261.9 404.6 349.3 137.4 41.1 39.5 53.6 161.9
VQIR 122.7 198.3 185.1 188.2 199.4 263.0 182.7 334.5 219.4
MS-ILLM 50.7 71.2 71.7 -38.4 -4.8 -37.3 -22.4 -6.0 4.1
PerCo 104.0 104.7 98.9 524.5 216.8 433.8 171.5 188.6 237.2
DiffEIC 21.9 22.4 28.3 514.6 68.9 158.1 48.7 48.5 108.2
RDEIC (Ours) 0 0 0 0 0 0 0 0 0
4.2 Experimental Results
Quantitative comparisons. Fig. 4 shows the performance of the proposed and compared methods on the CLIC2020 dataset. It can be observed that the proposed RDEIC demonstrates superior performance across different perceptual metrics compared to other methods, particularly achieving optimal results in DISTS, FID, and KID. For the distortion metrics, RDEIC significantly outperforms other diffusion-based methods, underscoring its superiority in maintaining consistency. To provide a more intuitive comparison of overall performance, we set our method as the anchor and compute the BD-rate (Bjontegaard, 2001) for each metric. The results are shown in Table 1. We observe that our method outperforms all the comparison methods, where the average BD-rate values of the competing methods are positive. In addition, the comparison results on the Tecnick and Kodak datasets are shown in Fig. 9 and Fig. 10, respectively, in Appendix C.
Qualitative comparisons. Fig. 5 provides visual comparisons among the evaluated methods at extremely low bitrates. MS-ILLM (Muckley et al., 2023) excels at reconstructing structural information, such as text, but falls significantly short in preserving textures and fine details. Diffusion-based Text+Sketch (Lei et al., 2023), PerCo (Careil et al., 2024) and DiffEIC (Li et al., 2024b) achieve realistic reconstruction at extremely low bitrates but often generate details and structures that are inconsistent with the original image. In comparison, the proposed RDEIC produces reconstructions with higher visual quality, fewer artifacts, and more faithful details. More visualization results can be found in Fig. 11 and Fig. 12 in Appendix C.
Complexity comparisons. Table 2 summarizes the average encoding/decoding times along with standard deviations for different methods on the Kodak dataset. Similar to PerCo (Careil et al., 2024), our RDEIC uses different denoising steps for different compression ratios. Due to the reliance on an iterative denoising process, diffusion-based extreme image compression methods require more time to reconstruct images than the GAN-based MS-ILLM (Muckley et al., 2023) and VQIR (Wei et al., 2024). Benefiting from the proposed relay residual diffusion, our RDEIC achieves high-quality image reconstruction with fewer denoising steps compared to other diffusion-based methods, significantly reducing the decoding time. For instance, compared to DiffEIC (Li et al., 2024b), our RDEIC is about 7 times faster using 5 denoising steps, and this speed advantage increases to a factor of 12 when using 2 denoising steps.
@@ 4.2 Experimental Results
Original MS-ILLM Text+Sketch PerCo DiffEIC RDEIC (Ours)
@@
Figure 5: Visual comparisons of our method to baselines on the CLIC2020 dataset. The bpp and DISTS of each method are shown at the bottom of each image. Compared to other methods, our method produces more realistic and faithful reconstructions.
Table 2: Encoding and decoding time (in seconds) on Kodak dataset. BD-rate (%) is calculated on CLIC2020 dataset, with DISTS as the metric. DS denotes denoising steps.
Types Methods DS Encoding Time Decoding Time BD-rate (%)
GAN-based MS-ILLM $0.038$$\pm$ $0.004$ $0.059$$\pm$ $0.004$ 50.7
VQIR $0.050$$\pm$ $0.003$ $0.179$$\pm$ $0.005$ 122.7
Diffusion-based Text+Sketch 25 $62.045$$\pm$ $0.516$ $12.028$$\pm$ $0.413$ 12584.9
PerCo 5 $0.236$$\pm$ $0.040$ $0.778$$\pm$ $0.007$ 104.0
PerCo 20 $0.236$$\pm$ $0.040$ $2.670$$\pm$ $0.008$ 104.0
DiffEIC 50 $0.128$$\pm$ $0.005$ $4.574$$\pm$ $0.006$ 21.9
RDEIC (Ours) 2 $0.119$$\pm$ $0.003$ $0.379$$\pm$ $0050$
RDEIC (Ours) 5 $0.119$$\pm$ $0.003$ $0.643$$\pm$ $0040$

### v2

@@ III-B Relay Residual Diffusion
where we set $\lambda=\frac{\sqrt{\bar{\alpha}_{N}}}{\sqrt{1-\bar{\alpha}_{N}}}$ to ensure $\eta_{N}=1$. Substituting Eq. (8) into Eq. (2), the diffusion process can be further written as follows:
$\boldsymbol{z}_{n}$ $=\sqrt{\bar{\alpha}_{n}}(\boldsymbol{z}_{0}+\lambda\frac{\sqrt{1-\bar{\alpha}_{n}}}{\sqrt{\bar{\alpha}_{n}}}\boldsymbol{e})+\sqrt{1-\bar{\alpha}_{n}}\epsilon_{n},$ (9)
$=\sqrt{\bar{\alpha}_{n}}\boldsymbol{z}_{0}+\sqrt{1-\bar{\alpha}_{n}}\underbrace{(\lambda\boldsymbol{e}+\epsilon_{n})}_{\tilde{\epsilon}_{n}}.$ (10)
It is evident that Eq. (10) has the same structure as Eq. (1), allowing us to seamlessly incorporate stable diffusion into our framework. For the denoising process, the noise estimator $\epsilon_{\theta}$ is learned to predict $\tilde{\epsilon}_{n}$ at each time-step $n$. The optimization of noise estimator $\epsilon_{\theta}$ is defined as
$\mathcal{L}_{ne}$ $=\mathbb{E}_{\boldsymbol{z}_{0},\boldsymbol{z}_{c},\boldsymbol{c},n,\epsilon_{n}}\|\boldsymbol{z}_{0}-\hat{\boldsymbol{z}}_{0}\|^{2}_{2},$ (11)
$=\omega_{n}\mathbb{E}_{\boldsymbol{z}_{0},\boldsymbol{z}_{c},\boldsymbol{c},n,\epsilon_{n}}\|\tilde{\epsilon}_{n}-\epsilon_{\theta}(\boldsymbol{z}_{n},\boldsymbol{c},n)\|^{2}_{2},$ (12)
where $\omega_{n}=\frac{1-\bar{\alpha}_{n}}{\bar{\alpha}_{n}}$. After that, we can start from the compressed latent features $\boldsymbol{z}_{c}$ and reconstruct the image using Eq. (5) without knowing the residual $\boldsymbol{e}$.
Figure 3: The two-stage training strategy of RDEIC. (a) Independent training: we randomly pick a time-step $n$ and train each time-step $n$ independently. This ensures that the model effectively learns to remove added noise and residuals at every step. (b) Fixed-step fine-tuning: $L$ fixed denoising steps are used to iteratively reconstruct a noise-free latent features $\hat{\boldsymbol{z}}_{0}$ from $\boldsymbol{z}_{N}$, which is consistent with the inference phase.
III-C Fixed-Step Fine-Tuning Strategy
Most existing diffusion-based image compression methods adopt the same training strategy as DDPM [13], where each time-step is trained independently. However, the lack of coordination among time-steps can lead to error accumulation and suboptimal reconstruction quality. To address this issue, we employ a two-stage training strategy. As shown in Fig. 3(a), we first train each time-step $n$ independently, allowing the model to learn to remove noise and residuals at each step. The optimization objective consists of the rate-distortion loss, codebook loss [41] and noise estimation loss:
$\begin{split}\mathcal{L}_{stage\ I}=\underbrace{\|sg(\boldsymbol{l}_{p})-\hat{\boldsymbol{l}}_{p}\|_{2}^{2}+\beta\|sg(\hat{\boldsymbol{l}}_{p})-\boldsymbol{l}_{p}\|_{2}^{2}}_{codebook\ loss\ \mathcal{L}_{cb}}+\\ \underbrace{\lambda_{r}\|\boldsymbol{z}_{0}-\boldsymbol{z}_{c}\|_{2}^{2}+R(\hat{\boldsymbol{y}})}_{rate-distortion\ loss\ \mathcal{L}_{rd}}+\lambda_{r}\mathcal{L}_{ne},\end{split}$ (13)
@@ III-C Fixed-Step Fine-Tuning Strategy
To this end, we further employ a fixed-step fine-tuning strategy to eliminate the discrepancy between the training and inference phases. As shown in Fig. 3(b), in each training step, we utilize spaced DDPM sampling [42] with $L$ fixed time-steps to reconstruct the noise-free latent features $\hat{\boldsymbol{z}}_{0}$ from the starting point $\boldsymbol{z_{N}}$ and map $\hat{\boldsymbol{z}}_{0}$ back to the pixel space $\hat{\boldsymbol{x}}=\mathcal{D}(\hat{\boldsymbol{z}}_{0})$. The loss function used in this stage is as follows:
$\begin{split}L_{stage\ II}=\lambda_{r}\Bigl{(}\|\boldsymbol{x}-\hat{\boldsymbol{x}}\|_{2}^{2}+\lambda_{lpips}\mathcal{L}_{lpips}(\boldsymbol{x},\hat{\boldsymbol{x}})\Bigr{)}+\\ \mathcal{L}_{rd}+\mathcal{L}_{cb}+\lambda_{r}\|\boldsymbol{z}_{0}-\hat{\boldsymbol{z}}_{0}\|^{2}_{2},\end{split}$ (14)
@@
III-D Controllable Detail Generation
@@
Although the fixed-step fine-tuning strategy significantly improves reconstruction quality, it requires a fixed number of denoising steps in the inference phase, making it impossible to achieve a trade-off between smoothness and sharpness by adjusting the number of denoising steps [16]. To address this limitation, we introduce a controllable detail generation method that allows us to dynamically balance smoothness and sharpness without being constrained by the fixed-step requirement, which enables more versatile and user-specific image reconstructions.
@@
Since the compressed latent feature already contains image information, directly using stable diffusion’s noise estimator $\epsilon_{sd}$ to predict noise $\epsilon_{sd}(\boldsymbol{z}_{n},n)$ results in low-frequency reconstructed images, as shown in the second column of Fig. 10. Inspired by classifier-free guidance [43], we decompose the predicted noise $\epsilon_{\theta}(\boldsymbol{z}_{n},\boldsymbol{c},n)$ into a low-frequency control component $\epsilon_{sd}(\boldsymbol{z}_{n},n)$ and a high-frequency control component $\epsilon_{\theta}(\boldsymbol{z}_{n},\boldsymbol{c},n)-\epsilon_{sd}(\boldsymbol{z}_{n},n)$, and control the balance between smoothness and sharpness by adjusting the intensity of the high-frequency control component:
$\hat{\epsilon}_{n}=\epsilon_{sd}(\boldsymbol{z}_{n},n)+\lambda_{s}(\epsilon_{\theta}(\boldsymbol{z}_{n},\boldsymbol{c},n)-\epsilon_{sd}(\boldsymbol{z}_{n},n)),$ (15)
@@
where $\lambda_{s}$ is the guidance scale. By adjusting the value of $\lambda_{s}$, we can regulate the amount of high-frequency details introduced into the reconstructed image. In the following experiments, we set $\lambda_{s}=1$ by default unless otherwise specified.
IV Experiments
Figure 4: Quantitative comparisons with state-of-the-art methods in terms of perceptual metrics (DISTS $\downarrow$ / NIQE $\downarrow$ / LPIPS $\downarrow$ / FID $\downarrow$ / KID $\downarrow$ ) on the Kodak, Tecnick ,and CLIC2020 datasets. Solid lines are used for diffusion-based methods, while dashed lines represent other methods. For RDEIC, we use 2 denoising steps for the two models with larger bpp and 5 steps for the remaining models.
Figure 5: Quantitative comparisons with state-of-the-art methods in terms of fidelity (PSNR $\uparrow$ / MS-SSIM $\uparrow$) on the Kodak, Tecnick ,and CLIC2020 datasets. Since Text+Sketch [14] has ignored the pixel-level fidelity of the reconstruction results, we do not report its rate-distortion performance.
IV-A Experimental Setup
IV-A1 Datasets
@@
The proposed RDEIC is trained on the LSDIR [44] dataset, which contains 84,911 high-quality images. For evaluation, we use three common benchmark datasets, i.e., the Kodak [45] dataset with 24 natural images of 768$\times$512 pixels, the Tecnick [46] dataset with 140 images of 1200$\times$1200 pixels, and the CLIC2020 [47] dataset with 428 high-quality images. For the Tecnick and CLIC2020 datasets, we resize the images so that the shorter dimension is equal to 768 and then center-crop them with 768$\times$768 spatial resolution [11].
IV-A2 Implementation details
We use Stable Diffusion 2.1-base [^1] https://huggingface.co/stabilityai/stable-diffusion-2-1-base as the specific implementation of stable diffusion. Throughout all our experiments, the weights of stable diffusion remain frozen. Similar to DiffEIC [16], the control module in our RDEIC has the same encoder and middle block architecture as stable diffusion and reduces the channel number to 20% of the original. The variance sequence $\{\beta_{t}\}_{t=1}^{T}$ used for adding noise is identical to that in Stable Diffusion. The number $N$ of denoising steps is set to 300. The size of codebook is set to 16384.
@@ IV-A2 Implementation details
For training, we use the Adam [48] optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$ for a total of 300K iterations. To achieve different compression ratios, we train five models with $\lambda_{r}$ selected from {2, 1, 0.5, 0.25, 0.1}. The batch size is set to 4. As described in Section III-C, the training process is divided into two stages. 1) Independent training. During this stage, the initial learning rate is set to 1×$\times 10^{-4}$ and images are randomly cropped to 512$\times$512 patches. We first train the proposed RDEIC with $\lambda_{r}=2$ for 100K iterations. The learning rate is then reduced to 2×$\times 10^{-5}$ and the model is trained with target $\lambda_{r}$ for another 100K iterations. 2) Fixed-step fine-tuning. In this stage, the learning rate is set to 2×$\times 10^{-5}$ and images are randomly cropped to 256$\times$256 patches. We fine-tune the model through the entire reconstruction process for 100K iterations. When $\lambda_{r}\in\{2,1\}$, the fixed number $L$ is set to 2, otherwise, it is 5. All experiments are conducted on a single NVIDIA GeForce RTX 4090 GPU.
IV-A3 Metrics
@@
For quantitative evaluation, we employ several established metrics to measure the visual quality of the reconstructed images, including reference perceptual metrics LPIPS [49], DISTS [50], FID [51] and KID [52] and no-reference perceptual metric NIQE [53]. We also employ distortion metrics PSNR and MS-SSIM [54] to measure the fidelity of reconstructions. Note that FID and KID are calculated on 256$\times$256 patches according to [9]. Since the Kodak dataset is too small to reliably calculate FID and KID scores, we do not report these results for this dataset.
IV-B Comparisons With State-of-the-art Methods
@@
We compare the proposed RDEIC with several representative extreme image compression methods, including the traditional standards: BPG [2] and VVC [3]; VAE-based method: ELIC [20]; GANs-based methods: HiFiC [9], MS-ILLM [10], and VQIR [29]; and diffusion-based methods: Text+Sketch [14], PerCo [15],and DiffEIC [16].
@@
Original VVC MS-ILLM Text+Sketch PerCo DiffEIC RDEIC (Ours)
@@
Figure 6: Visual comparisons of our method to baselines on the CLIC2020 dataset. Compared to other methods, our method produces more realistic and faithful reconstructions.
TABLE I: Encoding and decoding time (in seconds) on Kodak dataset. Decoding time is divided into the time spent in the denoising stage and the time spent in the remaining parts. The testing platform is RTX4090.
Types Methods Denoising Step Encoding Time Decoding time
Denoising Time Remaining Time
VAE-based ELIC 0.056 $\pm$ 0.006   – 0.081 $\pm$ 0.011
GAN-based HiFiC 0.038 $\pm$ 0.004   – 0.059 $\pm$ 0.004
MS-ILLM 0.038 $\pm$ 0.004   – 0.059 $\pm$ 0.004
VQIR 0.050 $\pm$ 0.003   – 0.179 $\pm$ 0.005
Diffusion-based Text+Sketch 25 62.045 $\pm$ 0.516 8.483 $\pm$ 0.344 4.030 $\pm$ 0.469
DiffEIC 50 0.128 $\pm$ 0.005 4.342 $\pm$ 0.013 0.228 $\pm$ 0.026
PerCo 5 0.236 $\pm$ 0.040 0.623 $\pm$ 0.003 0.186 $\pm$ 0.002
20 0.236 $\pm$ 0.040 2.495 $\pm$ 0.009 0.186 $\pm$ 0.002
RDEIC (Ours) 2 0.119 $\pm$ 0.003 0.173 $\pm$ 0.001 0.198 $\pm$ 0.003
5 0.119 $\pm$ 0.003 0.434 $\pm$ 0.002 0.198 $\pm$ 0.003
IV-B1 Quantitative Comparisons
Fig. 4 shows the rate-perception curves of the proposed and compared methods across the three datasets. It can be observed that the proposed RDEIC demonstrates superior performance across different perceptual metrics compared to other methods, particularly achieving optimal results in DISTS, FID, and KID. For the distortion metrics, as shown in Fig. 5, RDEIC significantly outperforms other diffusion-based methods, underscoring its superiority in maintaining consistency. Moreover, we report the performance of the SD autoencoder in Fig. 4 and Fig. 5 (see the black horizontal line), which represents the upper bound of RDEIC’s performance. Compared to DiffEIC [16], which is also based on stable diffusion, RDEIC is significantly closer to this performance upper limit.
IV-B2 Qualitative Comparisons
Fig. 1 and Fig. 6 provides visual comparisons among the evaluated methods at extremely low bitrates. VVC [3] and MS-ILLM [10] excel at reconstructing structural information, such as text, but falls significantly short in preserving textures and fine details. Diffusion-based Text+Sketch [14], PerCo [15] and DiffEIC [16] achieve realistic reconstruction at extremely low bitrates but often generate details and structures that are inconsistent with the original image. In comparison, the proposed RDEIC produces reconstructions with higher visual quality, fewer artifacts, and more faithful details.
IV-B3 Complexity Comparisons
Table I summarizes the average encoding/decoding times along with standard deviations for different methods on the Kodak dataset. For diffusion-based methods, decoding time is divided into denoising time and remaining time. Due to relay on stable diffusion, diffusion-based extreme image compression methods have higher encoding and decoding complexity than other learned-based methods. By reducing the number of denoising steps required for reconstruction, the denoising time of RDEIC is significantly lower than that of other diffusion-based methods. For instance, compared to DiffEIC [16], our RDEIC is approximately 10$\times$ to 25$\times$ faster in terms of denoising time.